OpenStudy (anonymous):

Which equation represents the line that contains the point (-2, 9) and is perpendicular to the line y = 2x + 6? A. y = -(1/2)x – 7 B. y = -(1/2)x + 8 C. y = -(1/2)x + 11 D. y = 2x + 13

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

@aroub

OpenStudy (aroub):

Yes, I can help you with that.

OpenStudy (aroub):

y=mx+b; m=-1/2, (-2,9) => 9=-1/2(-2)+b 9=1+b 8=b => y=-1/2x+8